OpenStudy (anonymous):

What is the average of all of the integers from 13 to 37?

OpenStudy (amistre64):

the midpoint between 13 and 37 of course

OpenStudy (amistre64):

add them then divide in half

OpenStudy (amistre64):

\[\sum_{13}^{37}=\frac{n(13+37)}{2}\] \[avg\sum_{13}^{37}=\frac{n(13+37)}{2}\div n=\frac{13+37}{2}\]